# Comprehensive Assessment of Jailbreak Attacks Against LLMs

### v2

@@ Abstract
Jailbreak attacks aim to bypass the safeguards of LLMs. While researchers have studied different jailbreak attacks in depth, they have done so in isolation—either with unaligned experiment settings or comparing a limited range of methods. To fill this gap, we present the first large-scale measurement of various jailbreak attack methods. We collect 17 cutting-edge jailbreak methods, summarize their features, and establish a novel jailbreak attack taxonomy. Based on eight popular censored LLMs and 160 questions from 16 violation categories, we conduct a unified and impartial assessment of attack effectiveness as well as a comprehensive ablation study. Our extensive experimental results demonstrate that all the jailbreak attacks have a powerful effect on the LLMs. This indicates that all LLMs fail to cover all the violation categories, and they are susceptible to significant jailbreak risks, with even the well-aligned Llama3 facing a maximum attack success rate of 0.88. Additionally, we test jailbreak attacks under eight advanced external defenses and find none of the defenses could mitigate the jailbreak attacks entirely. Our study offers valuable insights for future research on jailbreak attacks and defenses and serves as a benchmark tool for researchers and practitioners to evaluate them effectively.
@@ Introduction
Large Language Models (LLMs) have demonstrated outstanding capabilities in different domains. Nevertheless, such LLMs have raised concerns about being misused [80, 44, 37, 65] due to their amazing potential. In response, various regulations from different governments or organizations [1, 2, 3, 4] have emerged. These regulations aim to oversee the development and deployment of LLMs and limit the risk of misuse. Meanwhile, LLM-related service providers, such as Meta [5] and OpenAI [6], are constantly updating their usage policies and are incorporating methods such as reinforcement learning from human feedback (RLHF) to ensure LLMs’ alignment with human values [69, 60].
However, people have still found various methods to bypass the security mechanisms of LLMs and induce them to produce content that violates usage policies. It started with discovering that certain prompts could lead LLMs to generate inappropriate content, such as the popular example of instructing ChatGPT with “Do Anything Now (DAN).” [^1] https://www.washingtonpost.com/technology/2023/02/14/chatgpt-dan-jailbreak/. Subsequently, more and more prompts of this nature have been curated, either intentionally or unintentionally, which are referred to as jailbreak prompts in the wild [66]. Parallel with jailbreak prompts in the wild, researchers have also been exploring methods for automatically generating jailbreak prompts, such as GCG [81] and AutoDAN [49]. As shown in Figure 1, these methods have attempted to craft jailbreak prompts to bypass LLM safeguards. Research [40] even suggests that high success rate jailbreak attacks can be carried out on well-aligned LLMs without jailbreak prompts simply by exploiting the generation settings.
Figure 1: Examples of different jailbreak attack methods.
Despite the endlessly emerging jailbreak methods, a unified, systematic, and comprehensive fair benchmark for them is currently lacking. In particular, previous jailbreak attacks [54, 31] often compare with a limited set of jailbreak methods, and the experimental setups do not ensure alignment. Additionally, the continuing updates to LLMs’ built-in safeguards and the development of external defenses can render many jailbreak attacks unsuccessful. Therefore, it is also crucial to study how the performance of jailbreak methods changes in response to model updates or external defenses.
Our Contribution. We fill this gap by conducting the first unified comprehensive assessment of various jailbreak attack methods far beyond previous “jailbreak prompts in the wild.”
As illustrated in Figure 2, we collect 17 state-of-the-art jailbreak attacks and establish a novel taxonomy. The taxonomy is based on whether the original question undergoes alteration (such as translation, encoding, or adjustment with prefixes/suffixes) and, if altered, how these modified prompts are produced. This taxonomy contains six categories: human-based, obfuscation-based, heuristic-based, feedback-based, fine-tuning-based and parameter-based method. We summarize these features, detailing their access to the LLMs and requirements of the initial seeds. We use the results of querying the target LLM directly without jailbreaking as the baseline. These provide us with a comprehensive spectrum of jailbreak attacks against LLMs.
Figure 2: Overview of our measurement process.
We further construct a comprehensive diverse forbidden question set, tagging questions into 16 violation categories of our unified policy derived from five leading LLM-related service providers’ usage policies: Google [7], OpenAI [6], Meta [5], Amazon [18, 19], and Microsoft [20, 21].
We systematically measure the efficacy of various jailbreak methods across eight target LLMs and conduct a comprehensive ablation study, which includes attack time efficiency, prompt token length, transferability, and attack performance over time. We also evaluate these attacks under eight advanced jailbreak defenses.
Our findings indicate that none of the tested LLMs could completely avoid all violations, with even the most secure model, Llama2, exhibiting a baseline attack success rate of 0.31. We observe that all LLMs are susceptible to significant jailbreak risks, with even the well-aligned Llama3 and Llama3.1 facing a maximum attack success rate of 0.88 and 0.55, respectively. We identify the most effective jailbreak method, LAA [24], which achieved an average ASR of 0.85 across all tested LLMs and notably achieved a perfect ASR of 1.00 on both Vicuna and GPT-3.5. Our results also show that no defense can perfectly defend against all tested attacks.
Based on our taxonomy, we identify commonalities among different methods and reveal valuable insights. For instance, the heuristic-based methods usually have good attack performance and transferability. However, they are completely defended by even lightweight defense models, as they require initial seeds, and the generated jailbreak prompts have similar patterns.
Overall, we make the following key contributions.
  • We propose a novel jailbreak attack taxonomy to summarize jailbreak attacks into six categories and provide a high-quality, diverse forbidden question dataset covering 16 violation types from a unified usage policy.
  • We perform the first holistic unified evaluation of jailbreak attacks and reveal that all LLMs are at risk of jailbreaking. Additionally, our ablation studies shed light on the unique characteristics of different jailbreak attacks.
  • We evaluate these jailbreak attacks under eight advanced defenses. The results show that no defense can perfectly defend against all tested attacks, and jailbreak attacks requiring initial seeds are not robust to defenses.
  • Our findings, based on the taxonomy, reveal commonalities among different methods, providing valuable insights beneficial for developing future jailbreak attacks and defenses.
Background
@@
Safety training for LLMs is of utmost importance. These models possess a remarkable aptitude for understanding external information, such as in-context learning [55], and their proficiency in utilizing search engines like Bing with Copilot. [^2] https://copilot.microsoft.com/. However, the abundance of training data exposes LLMs to the risk of obtaining and distributing potentially harmful or unsafe knowledge. Adversaries exploit these capabilities to launch a variety of attacks, notably prompt injection [22] and jailbreak attacks [66, 34, 31, 49, 40]. To defend against these risks, LLMs have been trained in many safety guard techniques, including reinforcement learning from human feedback (RLHF) [27, 25] and red teaming [63].
@@
Currently, the majority of jailbreak attacks are accomplished through the creation of “jailbreak prompts.” These prompts are specialized inputs that exploit potential loopholes or weaknesses in the LLMs. Researchers have proposed various approaches for identifying or crafting jailbreak prompts, including collecting them from real-world scenarios [66], manually creating them by guided strategies [76, 72], or automatic generation [54, 33, 77, 31]. The previous work [40] also found that the alignment of LLMs cannot cover all hyperparameters, generating harmful content under specific hyperparameters without requiring the prompts.
On the other hand, defenses against jailbreak attacks have been developed to protect the LLMs using different perspectives. Some previous works [23, 42] exploit the high perplexity of jailbreak prompts for detection, while others [53] rely on pre-trained classifiers. Recently, some advanced works [46, 41] have employed another LLM to help detect and identify jailbreak prompts.
@@
Table 1: Summarization of jailbreak attack taxonomy. It details the access required, the necessity for modifications to the original question, and whether specific jailbreak prompts are utilized as initial jailbreak seeds.
@@
Parameter- Based Generation Exploitation /
Here, we present our attack taxonomy. We collect 17 advanced jailbreak methods, most of which are open-source or available online. We classify them based on the following two criteria:
  • (1)
    We first examine whether the original forbidden questions are altered to circumvent the target LLM’s alignment mechanisms within the method.
  • (2)
    Should the original question be altered, we then analyze the techniques used to generate these modified prompts in the method, such as by employing translations or by adding prefixes and suffixes.
We find that the 17 jailbreak methods could be summarized into six categories, i.e., human-based, obfuscation-based, heuristic-based, feedback-based, fine-tuning-based, and parameter-based method. The first five categories modify the original forbidden question to execute attacks, but their prompt generation methods differ significantly, while the sixth category solely uses the original question. We believe our attack taxonomy could cover most current jailbreak attacks. We summarize the features of each jailbreak method in Table 1.
@@
This category refers to jailbreak prompts generated by human-based method, e.g., the jailbreak prompts we use in the paper are collected from the contributors on the Internet. [^3] https://www.jailbreakchat.com. In the previous work [66], these prompts are also termed “jailbreak prompts in the wild.” These jailbreak prompts require no alteration to achieve the attack goal. In this scenario, the adversary is assumed to have black-box access to the target LLMs. In our paper, we select the top three human-based jailbreak prompt sets in “Votes” from the above website, including AIM, Devmoderanti, and Devmode v2.
@@
This category is the obfuscation-based method, which is a systematic and intentional approach that uses some obfuscation or non-English translation to jailbreak the LLMs. Such methods exploit vulnerabilities in the alignment mechanism. The adversary is assumed to have black-box access to the LLMs. Here, we select four common jailbreak methods.
Base64 [72]. Many LLMs [8, 59, 9] can recognize the Base64 encoding and thus the adversary could obfuscate the forbidden questions through Base64 encoding to bypass the safety mechanisms of LLMs.
@@
Methods in this category automatically optimize the jailbreak prompts with different heuristic optimization algorithms [79, 62], including mutation, random search, and genetic algorithm. Heuristic-based algorithms typically exhibit greater complexity, necessitating using specific human-crafted jailbreak prompts as initial seeds to reduce the search space.
AutoDAN [49]. AutoDAN automatically generates stealthy jailbreak prompts by modifying the initial seeds with a carefully designed hierarchical genetic algorithm. The adversary is assumed to have the white-box access to the LLMs.
@@
Methods in this category modify jailbreak prompts in a targeted manner based on feedback received during iterations, such as gradient information or jailbreak scores. Being less complex, they consequently rely less on human-based jailbreak prompts as the initial seed.
COLD [36]. This technique adapts Langevin dynamics [73] to perform efficient gradient-based sampling in the continuous logit space to conduct attacks. The adversary is assumed to have the white-box access to the LLMs.
@@
In this category, the adversary is required to fine-tune an LLM using the jailbreak prompts dataset as their attack model. Although the fine-tuning process is time-consuming, once it is completed, jailbreak prompts can be generated rapidly.
MasterKey [33]. MasterKey fine-tunes an LLM on various successful jailbreak prompts to learn effective patterns. Then the fine-tuned LLM could rewrite the input human-based jailbreak prompts (which may be invalid) to generate successful ones. Due to the unavailable source code, we rewrite AIM with the top-1 jailbreak template in their paper. The adversary is assumed to have black-box access.
Parameter-Based Method
Methods in this category manage to jailbreak the target LLM by exploiting the sampling methods or parameters during the generation process without creating typical jailbreak prompts.
Generation Exploitation [40]. It is an approach that disrupts model alignment by only manipulating the generation hyperparameters or variations of decoding methods. The adversary is assumed to have the white-box access to the LLMs.
@@
Policy Unification. LLM-related service providers are rapidly revising their usage policies to incorporate more privacy and safety concerns. However, these policies still exhibit notable variations among different providers. Therefore, to effectively evaluate all the jailbreak methods, it is paramount to formulate a unified policy covering safety questions across different LLMs. This unified policy acts as the foundational guideline, under which we gather forbidden questions from various categories to create a comprehensive test dataset.
@@
We first collect the usage policies from five major LLM-related service providers (Google [7], OpenAI [6], Meta [5], Amazon [18, 19], and Microsoft [20, 21]). Many policies tend to provide a general description by synthesizing many specific categories within an overarching category. Unlike the general ones, we summarize our unified policy by explicit coverage to find a clear common feature within the same category. Based on previous work [41], we categorize the usage policy into 16 violation categories (see Table 8 in Section A.1). We list the categories explicitly included in the policy of each LLM-related service provider in Table 9 in Section A.1.
@@
Dataset Establishment. We identify some redundancies in prior forbidden question datasets; for example, AdvBench [81] contains 24 bomb-related queries. And some strictly forbidden questions—like those about Crimes Involving Children—are included in previous works [81]. To address these, we initially handpick questions from prior works [81, 66], followed by a filtering process to remove improper, duplicates, or irrelevant queries based on our policy. To ensure the diversity and comprehensiveness of our dataset, we also employ the method in [66] with a designed prompt (refer to Figure 7 in Section A.4) to generate additional forbidden questions, which are then manually screened.
Overall, the forbidden question dataset is composed of 160 forbidden questions (ten questions for each violation category) with high diversity. Two human annotators manually review each question to ensure it indeed violates the corresponding category. Compared to previous works [81, 66], our dataset encompasses a wider range of categories and includes a more diverse array of questions. We use this dataset to query LLMs and serve as our baseline directly.
Note. Our criterion for determining whether a question belongs to a violation category relies on human annotation to ensure that the question classification aligns with human values and to eliminate potential machine generation biases. Questions related to Crimes Involving Children are strictly forbidden, so we exclude this topic from our dataset.
Experiment Settings
Target LLMs
We selected eight popular safe-aligned LLMs to assess the jailbreak attacks, covering both open-source LLMs, i.e., ChatGLM3 (chatglm3-6b[10], Llama2 (llama2-7b-chat[69], Llama3 (llama3-8b-instruct[11], Llama3.1 (llama3,1-8b-instruct[12], and Vicuna (vicuna-7b[13] and closed-source LLMs, i.e., GPT-3.5 (gpt-3.5-turbo[8], GPT-4 (gpt-4[59], PaLM2 (chat-bison@001[14], for different attack scenarios.
Datasets
We use the forbidden question dataset throughout the assessment of jailbreak attacks (see  Section 4). Some banned questions are manually selected from AdvBench [81] and Forbidden_Question_Set [66].
Evaluation Metrics
We adopt attack success rate (ASR) as our evaluation metric. ASR is the ratio of successful jailbreak queries $n$ to total queries $m$.
$\text{{ASR}}=\frac{n}{m}$
Here, we evaluate the top-1 attack success rate, wherein we generate a single response with the highest likelihood for each jailbreak candidate prompt and assess its effectiveness. More specifically, Generation Exploitation requires modification of hyperparameter settings and decoding methods, leading us to generate 50 response candidates for every forbidden question. If any of the responses is labeled as successful, the corresponding question is classified as a successful attack.
In addition, how to determine the success of jailbreak is still an open question, especially for large queries. We conduct human annotation and summarize some common responses prone to be misclassified and their correct labels with explanations (see Table 11 in Section A.3). We introduce GPT-4 as our judging model to enhance the evaluation of the ASR. We utilized few-shot learning [29], whereby the judging model can label the responses with three demonstrations (see Figure 8 in Section A.4 for the judge prompt). Specifically, labels need to be assigned from three different perspectives:
To evaluate the effectiveness of our evaluation method, we manually check the classification results. We randomly select 640 responses, which are then independently labeled by three different annotators. We employ the majority vote to resolve inconsistencies in labeling. K-alpha value [45] of labeling is 0.87, indicating a good consistency among the three annotators. Additionally, the label matching rate is 94.84% of 640 responses, signifying that our proposed method establishes strong stability when compared to human-labeled results.
Hyperparameters
@@
Baseline. We use the forbidden questions without employing jailbreak techniques to query the target LLMs directly. The results in this setting serve as the baseline for our experiment.
@@
Unification of the Term “Step” During Optimization. Different jailbreak methods have varying definitions of the term “step.” For instance, GCG reports the number of optimizing epochs as the step, while TAP sets the total count of queries to the target LLMs as the step. Note that, for TAP, there are still some queries to the evaluator from the generated response candidates. Therefore, it is unfair to compare the steps defined in different jailbreak methods directly. To address this, we adopt a general definition of “step” in our experiments. Each modification of the prompts is regarded as one step in employing auxiliary LLMs to modify jailbreak prompts. The maximum number of modification steps for each forbidden question is set to 50. We refer to the step in GCG and COLD as “gcg_step” and set its number to 500. In this configuration, the performance and efficiency of GCG and COLD are comparable to those of other methods. For Generation Exploitation, within each generation, we select 50 different configurations of hyperparameters and decoding methods, resulting in 50 responses for each violated question.
@@
Remark. For each forbidden question, we generate a specific jailbreak candidate prompt using each jailbreak method on each target LLM, termed as “direct attack” in previous works [81, 49, 31]. We outline additional hyperparameters and experimental settings in Table 10 in  Section A.2.
Experiment Results

### v3

# JailbreakRadar: Comprehensive Assessment of Jailbreak Attacks Against LLMs

@@ Abstract
Jailbreak attacks aim to bypass the LLMs’ safeguards. While researchers have proposed different jailbreak attacks in depth, they have done so in isolation—either with unaligned settings or comparing a limited range of methods. To fill this gap, we present a large-scale evaluation of various jailbreak attacks. We collect 17 representative jailbreak attacks, summarize their features, and establish a novel jailbreak attack taxonomy. Then we conduct comprehensive measurement and ablation studies across nine aligned LLMs on 160 forbidden questions from 16 violation categories. Also, we test jailbreak attacks under eight advanced defenses. Based on our taxonomy and experiments, we identify some important patterns, such as heuristic-based attacks could achieve high attack success rates but are easy to mitigate by defenses, causing low practicality. Our study offers valuable insights for future research on jailbreak attacks and defenses. We hope our work could help the community avoid incremental work and serve as an effective benchmark tool for practitioners.
$\clubsuit$\$$\clubsuit$\$footnotetext: Corresponding author.
@@ Introduction
Large language models (LLMs) have garnered significant attention due to their exceptional capabilities across various domains. However, numerous instances of misuse [94, 37, 30, 72] have also emerged. In response, many governments and organizations have introduced corresponding regulations [24, 66, 2, 13], usage policies [49, 64], or safety measures [79, 67] to align LLMs with human values and mitigate the risk of misuse.
However, various jailbreak methods are being developed. They began with “jailbreak prompts in the wild,” referring to jailbreak prompts found online, such as the famous “Do Anything Now.” [^1] https://www.washingtonpost.com/technology/2023/02/14/chatgpt-dan-jailbreak/. Subsequently, researchers try to generate new jailbreak prompts automatically by mutating existing ones [42] or leveraging gradient information from LLMs [96]. Additionally, researchers [33] have found that, even without using jailbreak prompts, simply altering the inference parameters can bypass restrictions and jailbreak the LLMs. We show examples of different jailbreak attack settings in Figure 1.
Figure 1: Examples of different jailbreak settings.
Despite the endlessly emerging jailbreak methods, there lacks a unified, systematic, and comprehensive fair benchmark. Particularly, previous jailbreak attacks [48, 15] often compare with a limited set of jailbreak methods, and some of their experimental setups do not ensure alignment. Also, some previous evaluation works [74, 82, 73, 43] solely investigate human-based or obfuscation-based attacks, without including new automatic methods.
Our Contribution. We fill such gaps by conducting a unified holistic assessment of jailbreak attacks, the first covering multiple attack types, including both automatic and non-automatic ones. Additionally, we perform comprehensive ablation studies and evaluations under various defense mechanisms, providing insights beyond merely reporting attack success rates (ASRs).
Figure 2: Overview of our assessment process.
Our assessment pipeline is shown in Figure 2. We first collect 17 representative jailbreak attacks and establish a novel attack taxonomy. Specifically, the categorization in our taxonomy depends on whether the attack requires additional jailbreak prompts and how these jailbreak prompts are produced. This taxonomy contains six categories: human-based, obfuscation-based, heuristic-based, feedback-based, fine-tuning-based and generation-parameter-based method. We further construct a comprehensive, diverse forbidden question set, tagging questions into 16 violation categories of our unified policy derived from five leading LLM-related service providers’ usage policies [27, 64, 49, 4, 5, 55, 56]. Then, we systematically measure the efficacy of various jailbreak methods across nine LLMs and conduct comprehensive ablation studies. We also evaluate these attacks under eight advanced jailbreak defenses.
Main Findings. Based on our taxonomy and assessment, the main findings are outlined below:
  • In real-world black-box settings, even the latest LLMs face significant jailbreak risks. For example, LAA achieves a 100% ASR on DeepSeek-V3.
  • Although Human-based, heuristic-based and other attacks using initial seeds could achieve high ASRs, their jailbreak prompts lack diversity and exhibit similar distributions, making them vulnerable to defenses that render them nearly ineffective. For example, PromptGuard can reduce LAA’s attack success rate to 0%.
  • Methods that generate diverse and natural jailbreak prompts, such as most feedback-based attacks, exhibit more stable attack performance and are relatively less affected by defenses. For example, PAIR and TAP still achieve ASRs above 15% even when all eight defense strategies are deployed.
Implications. We hope the diverse forbidden question dataset we constructed—spanning 16 violation categories across five leading LLM providers (the most comprehensive to date)—to be reusable in future research. Moreover, we wish the insightful observations based on our taxonomy to help the community avoid incremental work, such as giving lower priority to heuristic-based attacks.
Background and Related Works
In this section, we mainly introduce related aligned LLMs and jailbreak attacks and defenses. We also discuss more related works in Appendix I, including the misuse and security measures of LLMs.
@@
Safety training for LLMs is of utmost importance. These models possess a remarkable aptitude for understanding external information, such as in-context learning [57], and their proficiency in utilizing search engines like Bing with Copilot. [^2] https://copilot.microsoft.com/. However, the abundance of training data exposes LLMs to the risk of obtaining and distributing potentially harmful or unsafe knowledge. Adversaries exploit these capabilities to launch a variety of attacks [1, 74, 22, 15, 42, 17, 33]. To defend against these risks, LLMs have been trained in many safety guard techniques, including reinforcement learning from human feedback (RLHF) [10, 8] and red teaming [70].
@@
Most jailbreak attacks are accomplished through the creation of “jailbreak prompts.” These prompts are specialized inputs that exploit potential loopholes or weaknesses in the LLMs. Researchers have proposed various approaches for collecting or crafting jailbreak prompts, including collecting them from real-world scenarios [74], manually creating them by guided strategies [88, 82], or automatic generation [48, 21, 89, 15]. The previous work [33] also found that the alignment of LLMs cannot cover all generation parameters, generating harmful content under specific parameters without altering the original questions.
Defenses against jailbreak have been developed to protect the LLMs using different perspectives. Some previous works [3, 35] exploit the high perplexity of jailbreak prompts for detection, while others [47] rely on pre-trained classifiers. Recently, some advanced works [39, 34] have employed another LLM to help detect and identify jailbreak prompts.
Previous evaluation works [74, 82, 73, 43] provide important insights into jailbreak but solely cover those non-automatic human-based or obfuscation-based attacks. Unlike theirs, our work includes both non-automatic and newly emerging automatic jailbreak attacks as well as comprehensive ablation studies. [^3] Within 12 months, we have several concurrent works. We discuss some of them in Appendix I.3.
@@
Table 1: Summarization of tested jailbreak attacks.
@@
Generation- Parameter-Based Generation Exploitation (GE) /
We collect 17 representative jailbreak attacks (details in Appendix G), and classify them based on two criteria: (C1) We first examine whether the original forbidden questions are altered to circumvent the target LLM’s alignment mechanisms within the method. (C2) Should the original question be altered, we then analyze the techniques used to generate these modified prompts in the method, such as by employing translations or by adding prefixes and suffixes.
Based on C1, we identify generation-parameter-based methods, which solely use the original questions. Based on C2, we further identify five other categories, including human-based, obfuscation-based, heuristic-based, feedback-based, fine-tuning-based. These five categories modify the original forbidden question to execute attacks (i.e., they require jailbreak prompts), but their prompt generation methods differ significantly. We believe our attack taxonomy could cover most current jailbreak attacks and summarize the features of each jailbreak method in Table 1.
Note. Our attack taxonomy mainly focuses on how the attacks jailbreak the target LLMs, instead of other features like “access.”
@@
Description. Human-based methods refer to those using “jailbreak prompts in the wild” [74], which are collected from the Internet.
Involved Attacks. AIM, Devmoderanti, and Devmode v2 (the top three human-based jailbreak prompts in “Votes” on “jailbreakchat” website). [^4] https://github.com/alexalbertt/jailbreakchat.
@@
Description. Obfuscation-based methods are those using some obfuscation (e.g., non-English translation) to jailbreak the LLMs. Such methods usually exploit vulnerabilities, such as low-resource languages or seemingly harmless synonyms, in the alignment mechanism.
Involved Attacks. Base64 [82, 73] (using Base64 coding), Combination [82] (using Base64, prefix&style injection), Zulu [88] (using low-resource language Zulu), and DrAttack [41] (using seemingly harmless synonyms).
@@
Description. Methods in this category automatically optimize the jailbreak prompts with different heuristic optimization algorithms [91, 69], including mutation, random search, and genetic algorithm. Heuristic-based algorithms typically necessitate using specific human-crafted jailbreak prompts as initial seeds to reduce the search space.
Involved Attacks. AutoDAN [42], GPTFuzz [89], and LAA [6].
@@
Description. Methods in this category modify jailbreak prompts in a targeted manner based on feedback received during iterations, such as gradient information or jailbreak scores. [^5] Drawing on the concepts from other domains, we take a broad definition of “feedback.” For example, in the field of automatic control, gradient descent-based learning is considered a feedback-based control strategy [76], as it relies on the loss between the response and the target to guide the next optimization step. Similarly, those methods that optimize the jailbreak prompts based on the jailbreak score and the objective also match the definition of feedback-based. Due to optimizing against the received feedback, these methods require less search and rely less on human-based jailbreak prompts as the initial seed.
Involved Attacks. GCG [96], COLD [29], PAIR [15], and TAP [48].
@@
Description. In this category, the adversary needs to fine-tune an attack LLM to conduct jailbreaks. The fine-tuned attack LLM could generate the potential jailbreak prompts according to the input forbidden questions.
Involved Attacks. MasterKey [21] and AdvPrompter [68].
Generation-Parameter-Based Method
Description. Methods in this category manage to jailbreak the target LLM by exploiting the sampling methods or parameters during the generation process without creating typical jailbreak prompts.
Involved Attacks. Generation Exploitation (GE) [33]. [^6] Generation-parameter-based attacks are relatively limited, but their mechanisms are fundamentally different from others. Thus, they need separate categorization and analysis.
@@
Policy Unification. LLM-related service providers are rapidly revising their usage policies to address more safety concerns. These policies also exhibit variations among different providers. Therefore, we aim to formulate a comprehensive unified policy covering safety concerns across different providers.
@@
We first collect the latest usage policies from five major LLM-related service providers (Google [27], OpenAI [64], Meta [49], Amazon [4, 5], and Microsoft [55, 56]). To the best of our knowledge, our study involves the largest number of providers. Many policies tend to provide a general description by synthesizing many specific categories within an overarching category. Unlike the general ones, we summarize our unified policy by explicit coverage to find a clear common feature within the same category. We then categorize the usage policy into 16 violation categories (see Table 8 in Appendix B). We list the categories explicitly included in the policy of each LLM-related service provider in Table 9 in Appendix B. We manually annotate 16 violation categories, classifying them into “general” (violations based on general human moral principles) and “specific” (violations that may be region-specific) to gain a deeper understanding of different violation categories (details in Section B.3).
@@
Dataset Establishment. We identify redundancies in prior datasets; for example, AdvBench [96] contains 24 bomb-related queries. And some strictly forbidden questions—like those about Child Endangerment—are included in previous works [96]. [^7] Detailed description in Appendix C.2. To address these, we first handpick questions from prior works [96, 74], followed by a filtering process to remove improper, duplicate, or irrelevant queries. To ensure the diversity and comprehensiveness of our dataset, we also employ the method in [74] with a designed prompt (refer to Figure 8 in Appendix J) to generate additional forbidden questions, which are then manually screened. Overall, the forbidden question dataset is composed of 160 forbidden questions (10 questions for each violation category) with high diversity. [^8] Child Endangerment is strictly forbidden, so we exclude it from our dataset (explanation details in Appendix B.2). Two human annotators manually review each question to ensure it indeed violates the corresponding category. Compared to previous works [96, 74], our dataset encompasses a wider range of categories and includes a more diverse array of questions.
Evaluation Settings
Test Datasets and LLMs. We use the forbidden question dataset built in Section 4. We select nine popular aligned LLMs. Five of them are in open-source settings, including ChatGLM3 (chatglm3-6b[93], Llama2 (llama2-7b-chat[79], Llama3 (llama3-8b-instruct[50], Llama3.1 (llama3.1-8b-instruct[51], and Vicuna (vicuna-7b[45]. Four of them are in closed-source settings, including GPT-3.5 (gpt-3.5-turbo[61], GPT-4 (gpt-4[63], DeepSeek-V3 (deepseek-v3-671b[20], PaLM2 (chat-bison@001[26]. [^9] DeepSeek-V3 is open-source, but due to the computing resource limitation, it is used under closed-source settings.
@@
Baseline. Directly querying the target LLMs using forbidden questions without jailbreak attacks serves as the baseline for our experiment.
Metrics. We adopt attack success rate (ASR) as our evaluation metric. ASR is the ratio of successful jailbreak queries $n$ to total queries $m$ ($\text{{ASR}}=\frac{n}{m}$).
How to determine the success of jailbreak on a large scale is also an open question. Previous studies have proposed string match [96] and LLM-as-a-judge [73, 92]. We conduct human annotation and find that previous methods are useful, but not ideal. Thus, we employ GPT-4-Turbo (gpt-4-turbo) as our judging model to label the responses from three aspects, aiming to evaluate the responses comprehensively and reduce misclassification. Our human annotation shows our method outperforms other methods (details in Appendix D.1).
@@
Unification of the Term “Step”. Different jailbreak methods, especially those automatic methods, have varying definitions of the term “step.” For instance, GCG reports the number of optimizing epochs as the step, while TAP sets the total count of queries to the target LLMs as the step. For TAP, there are still some queries to the evaluator from the generated response candidates. Therefore, it is unfair to compare the steps defined in different jailbreak methods directly. To address this, we adopt a general definition of “step” in our experiments. Each modification of the prompts is regarded as one step in employing auxiliary LLMs to modify jailbreak prompts. The maximum number of modification steps for each forbidden question is set to 50. We refer to the step in GCG and COLD as “gcg_step” and set its number to 500. In this configuration, the performance and efficiency of GCG and COLD are comparable to those of other methods. Note that some methods are not compatible with the concept of “steps,” as they are jailbreak attacks based on fixed templates rather than iterative processes. For example, Combination already represents the best-performing template and follows a fixed-format approach without involving iterative “steps.” Base64/Zulu are in the same situation.
Under these settings, we evaluate the top-1 ASR for all methods except GE, wherein we generate a single response with the highest likelihood for each jailbreak candidate prompt and assess its effectiveness. For GE, we select 50 different generation settings for each forbidden question, resulting in 50 responses. If any of the responses is labeled as successful, the corresponding question is considered a successful jailbreak.
@@
Remark. For each forbidden question, we conduct an individual attack using each jailbreak method on each target LLM, termed as “direct attack” in previous works [96, 42, 15]. Additional experimental settings are shown in Table 13 of  Appendix F.
Evaluation Results